OpenStudy (anonymous):

The prices of tickets for the local rock concert were $16 for Friday night and $28 for Saturday night. The total attendance for the two nights was 1090. If the total receipts were $24,232, how many people attended Saturday night?

OpenStudy (texaschic101):

F + S = 1090 ( total attendance for Friday and Saturday) 16F + 28S = 24,232 (16 for Friday and 28 for Saturday = 24,232) can you solve it from here or do you still need help ?

OpenStudy (anonymous):

uh do you divide or..

OpenStudy (texaschic101):

no...take your first equation F + S = 1090....and since we are looking for S, lets change the equation to : F = 1090 - S. Now sub 1090 - S in for F in the first equation. 16F + 28S = 24232 16(1090 - S) + 28S = 24232 -- distribute through the parenthesis 17440 - 16S + 28S = 24232 -- subtract 17440 from both sides -16S + 28S = 24232 - 17440 -- combine like terms 12S = 6792 -- divide both sides by 12 S = 6792/12 S = 566 566 peeps attended Saturday night

OpenStudy (anonymous):

okay, thank you

OpenStudy (texaschic101):

any questions at all ?

OpenStudy (anonymous):

no, thank you

OpenStudy (texaschic101):

cool..glad to help :)